OpenStudy (i_love_my_nieces):

help (m^2 - m + 3) + (m - 1) This is a polynomial

OpenStudy (i_love_my_nieces):

@Michele_Laino

OpenStudy (i_love_my_nieces):

@KendrickLamar2014

OpenStudy (i_love_my_nieces):

@jagr2713

OpenStudy (i_love_my_nieces):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Are you supposed to simplify it or something?

OpenStudy (kendricklamar2014):

If you have to Simplify it: \[m^2−m+3+m−1 = m^2+−m+3+m+−1\]

OpenStudy (kendricklamar2014):

Now you Combine-Like-Terms:\[m^2+−m+3+m+−1 = (m^2)+(−m+m)+(3+−1) = m^2+2\]

OpenStudy (kendricklamar2014):

\[Answer: m^2 + 2\]

OpenStudy (kendricklamar2014):

BTW: This is History not Mathematics :P

OpenStudy (i_love_my_nieces):

Oh oops